Let demand equal: P = 10 – 0.2Qd. Supply is denoted as: P = 2 + 0.2Qs. Assume that Qd is quantity demanded and Qs is quantity supplied.


1
Expert's answer
2021-03-05T11:19:11-0500

"solution"

"Given ;\\
demand \ equation:P=10-0.2Qd\\
supply\ equation \ :P=2+0.2Qs\\
equilibrium\ price\ and \ quantity \ can\ be\ determined \ as \ follows:\\"


"at\ equilibrium\ , \ Qd=Qs\\"


"10-0.2Q=2+0.2Q\\
10-2=0.2Q+0.2Q\\
8=0.4Q\\
Q=\frac{8}{0.4}\\
Q=20"


Quantity at equilibrium is "20" units.

The price at equilibrium is got by substituting "Q" in one of the equations :

"P=2+0.2Q\\
P=2+(0.2\times 20 )\\
P=2+4\\
P=6"


The price at equilibrium is "\$6" .